### v4

@@ 3.1 Convergence analysis
Theorem 11.
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
If $w(x_{i}^{*},0)<+\infty$, it follows from [48, Proposition 2.1.5], [42, Theorem 5.19] and the outer semi-continuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$
as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$. Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (6) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)<\infty,i\in\mathcal{J}^{*}$, (7) is satisfied by Assumption 2. Therefore, $x^{*}$ is stationary. ∎
3.2 Avoidance of strict saddle point
The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
Theorem 12.
(Model Identification and Local Smoothness) Under Assumption 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
@@ 3.2 Avoidance of strict saddle point
    $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
@@
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with following constrained problem
$S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (29)
@@
where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by
$g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
Therefore, $S$ is $C^{1}$ smooth in $B$.
(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$. ∎
@@
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (30)
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
Theorem 13.
(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$. Moreover, if $\alpha<\frac{\beta}{\rho}$, $DT(x^{*},\epsilon^{*})$ is invertible.
Proof.
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 12. It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P). Recall Definition 5 and Definition 3 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$. It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue. This equivalence is obviously true since
@@
If $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\lambda_{\max}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue, we know $DT(x^{*},\epsilon^{*})$ is invertible. ∎
To apply Corollary 1, we next show that $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
Proposition 14.
(Lipeomorphism) Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible; moreover, $T$ and $T^{-1}$ is Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
Proof.
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
$r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (31)
by Assumption 3. This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
with
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$. Let $L_{r}$ be the Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
$\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$ (32)
@@
Next we prove that $S$ is Lipschitz continuous. It follows that
@@
$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
@@
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
where the first inequality is determined by Lemma A.3 and the second is by (21) and (32). Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism. ∎
Now we are ready to apply Corollary 1 to derive the following global escape theorem.
Theorem 15.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3 satisfied. Suppose $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$, $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$ and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof.
It follows directly from Theorem 13 that the map $T$ is local $C^{1}$ smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points. By applying Proposition 14, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure. ∎
We illustrate this property by applying DIRL1 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points: $(3,3)$ representing an initial point far from the global minimizer and $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 3. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
(a) $\alpha=0.9$ and initial point $(3,3)$.
(b) $\alpha=0.9$ and initial point $(0,0.043)$.
(c) $\alpha=0.1$ and initial point $(3,3)$.
(d) $\alpha=0.1$ and initial point $(0,0.043)$.
Figure 3: Trajectories of DIRL1.

### v5

@@ 3.1 Convergence analysis
Theorem 3.5.
(Global convergence) Suppose Assumptions 1 and 3 hold. Let the sequence $\{(x^{k},\epsilon^{k})\}$ be generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is a stationary point of (P).
Proof 3.6.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem (26) implies $-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$ If $w(x_{i}^{*},0)<+\infty$, it follows from the outer semicontinuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that $0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$ as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$. Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (5) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)=\infty,i\in\mathcal{J}^{*}$, (6) is satisfied by Assumption 2. Therefore, $x^{*}$ is a stationary point.
3.2 Avoidance of strict saddle points
The map $S^{x}$ has the property of so-called “active manifold identification”, which means that the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [41, 42] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
Theorem 3.7.
(Model Identification and Local Smoothness) Under Assumptions 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
@@ 3.2 Avoidance of strict saddle points
    (28) $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
@@
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with the following constrained problem,
(29) $S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$
@@
where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by (28). Therefore, $S$ is $C^{1}$ smooth in $B$.
(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$.
@@
(30) $T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$
around $(x^{*};\epsilon^{*})$. We then proceed to the saddle escaping mechanism for DIRL1, ensuring that the algorithm does not converge to the strict saddle points defined in Definition 2.7. To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
Theorem 3.9.
(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$. Moreover, if $\alpha<\frac{\beta}{\rho}$ where $\rho$ is defined by (22), $DT(x^{*},\epsilon^{*})$ is invertible.
Proof 3.10.
First of all, we calculate the Jacobian of $T$ in the neighborhood $B$ described in Theorem 3.7. It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \boldsymbol{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \boldsymbol{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \boldsymbol{0}&(1-\alpha)I&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P). Recall Definition 2.12 and Definition 2.7 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$. It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue. This equivalence is obviously true since
@@
Moreover, if $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\rho>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have a 0 eigenvalue, we know by (22) that $DT(x^{*},\epsilon^{*})$ is invertible.
To apply Corollary 2.15, we next show that $T$ is a lipeomorphism, even if it may not be differentiable globally in $\mathbb{R}^{2n}$. This is equivalent to showing that $T$ and its inverse $T^{-1}$ are Lipschitz continuous.
Proposition 3.11.
(Lipeomorphism) Suppose Assumptions 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C)|)$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in (20), and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible; moreover, $T$ and $T^{-1}$ are Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{\nabla f}/2$ and $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$.
Proof 3.12.
If $r^{\prime}(|0^{+}|)<\infty$, $w(x_{i},\epsilon_{i})\leq r^{\prime}(|0^{+}|)$ by Assumption 1 and $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})$ is Lipschitz continuous for $x\in\mathcal{L}(F^{0})$ and $\epsilon_{i}\in\mathbb{R}_{+}$ by Assumption 2.
If $r^{\prime}(|0^{+}|)=\infty$, by (20), $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $w(x_{i},\epsilon_{i})>C/\lambda$, by the definition of $S^{x}$,
$0\in\nabla_{i}f(x)+\beta(y_{i}-x_{i})+\lambda w(x_{i},\epsilon_{i})\partial|y_{i}|.$
It is true that $S_{i}^{x}(x,\epsilon)=0$. Therefore, for any $x\in\mathcal{L}(F^{0})$ with $S_{i}^{x}(x,\epsilon)\neq 0$,
(31) $r^{\prime}(|x_{i}|+\epsilon_{i})\leq C/\lambda\text{ \ and \ }|x_{i}|+\epsilon_{i}\geq\underline{x}:=(r^{\prime})^{-1}(C/\lambda)>0$
by Assumption 1.
It implies that we can equivalently write $S$ as
$S_{i}^{x}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
where we define $u:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ with
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0^{+}|)<\infty,\\ \min(C/\lambda,w(x_{i},\epsilon_{i}))=\min(C/\lambda,r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0^{+}|)=\infty.\end{cases}$
Therefore, $u(\cdot)$ is also Lipschitz continuous for $x\in\mathcal{L}(F^{0})$ and $\epsilon\in\mathbb{R}^{n}_{+}$. Let $L_{r}$ be its Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
(32) $\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$
@@
Next, we prove that $S$ is Lipschitz continuous. It follows that
@@
$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}\|\hat{u}-\tilde{u}\|$
@@
$\leq$ $\ (1+\tfrac{1}{\beta}L_{\nabla f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
where the first inequality is determined by Lemma A.1 and the second is by (21) and (32).
Therefore, for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$,
$\|\alpha(I-S)(\hat{x},\hat{\epsilon})-\alpha(I-S)(\tilde{x},\tilde{\epsilon})\|$
$\leq$ $\ \alpha\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|+\alpha\|S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon})\|+\alpha\|S^{\epsilon}(\hat{x},\hat{\epsilon})-S^{\epsilon}(\tilde{x},\tilde{\epsilon})\|$
$\leq$ $\ \alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|+\alpha\mu\|\hat{\epsilon}-\tilde{\epsilon}\|$
$\leq$ $\ \alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$
Notice that $T=I+\alpha(S-I)$. Since $\alpha(I-S)$ is Lipschitz with constant $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$. Applying [15, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism.
Now we are ready to apply Corollary 2.15 to derive the following global escape theorem.
Theorem 3.13.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumptions 1-3 being satisfied. Suppose $10<\alpha<$ 1, $\beta>\alpha L_{\nabla f}/2$, $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$ and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 2. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof 3.14.
It follows directly from Theorem 3.9 that the map $T$ is local $C^{1}$ smooth around any stationary point, and that all stationary points are unstable fixed points. By applying Proposition 3.11, $T$ is a lipeomorphism. By Corollary 2.15, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure.